OpenStudy (anonymous):

The formula for the area of a triangle is A= 1/2bh, where b is the base of the triangle and h is the height of the triangle. What is the length of the base if the area is 22cm^2 and the height is 8cm?

OpenStudy (ajprincess):

A=1/2*b*h 2A=b*h b=(2A)/h b=(2*22)/8 Hope u can find b nw:)

OpenStudy (hba):

Be careful of the unit cm^2/cm=cm

OpenStudy (anonymous):

alright alright

OpenStudy (anonymous):

b=5.5cm

OpenStudy (ajprincess):

yup perfect:)

OpenStudy (anonymous):

thank you

OpenStudy (ajprincess):

welcome:)